Sorry my daughter does not have a grade 9 textbook...just handouts. The handouts are on 'word problems'. The heading on the handout is '5.4 Value Problems'
Brianne received 9 bills, some twenties,some tens, and some fives, in exchange for a $100 bill. There was twice as many fives as there were tens. How many of each kind of bill did she receive?
$5 = y= x(2)
$10 = x
$20 = z= 9-(x+y)

Answer by nerdybill(7384)

Brianne received 9 bills, some twenties,some tens, and some fives, in exchange for a $100 bill. There was twice as many fives as there were tens. How many of each kind of bill did she receive?
.
Let x = number of tens
then
2x = number of fives
9-(x+2x) = number of twenties
.
Since the total is $100:
10x + 5(2x) + 20(9-(x+2x)) = 100
10x + 5(2x) + 20(9-x-2x) = 100
10x + 5(2x) + 20(9-3x) = 100
10x + 10x + 180 - 60x = 100
180 - 40x = 100
-40x = -80
x = 2 (number of tens)
.
number of fives:
2x = 2(2) = 4
.
number of twenties:
9-(x+2x) = 9-3x = 9-3(2) = 9-6 = 3
.
Answer:
2 tens, 4 fives, and 3 twenties